# HOSIG: Full-Body Human-Object-Scene Interaction Generation with Hierarchical Scene Perception

### v1

@@ 3.4. Scene-Guided Controllable Motion Generation
Ours 0.027 13.0367 0.9333 0.1851 0.6562 201.8264 0.9800 0.0007 0.0113
@@
(9) $\mu_{t}=\mu_{t}-\tau\nabla_{\mu_{t}}\left(\lambda_{1}\mathcal{L}_{j}^{root}+\lambda_{2}\mathcal{L}_{j}^{hand}+\lambda_{3}\mathcal{L}_{s}\right)$
@@
4. Experiments
4.1. Implementation Details
@@
Our evaluation protocol comprehensively assesses both absolute performance and component effectiveness. All experiments are conducted on the TRUMANS dataset (Jiang et al., 2024b), containing 100 indoor scenes with annotated human-object interactions. We establish two evaluation axes: 1) comparison against SOTA methods in HOI and HSI generation, and 2) ablation studies isolating our core technical innovations.
SOTA Comparisons While no existing method achieves identical functionality, we select two representative baselines for fair comparison.
  • LINGO (Jiang et al., 2024a): A scene-conditioned motion generator. As their open-source implementation lacks object trajectory synthesis, we approximately reproduce the object interaction results using their paper’s methodology. Since it additionally requires predefined finger positions, we use the hand pose generated by our method as an additional condition.
  • CHOIS (Li et al., 2024a): A scene-aware motion model requiring pre-defined object paths. We feed CHOIS with trajectories from Algorithm 1 to ensure comparable navigation inputs. Note for fair comparison, we fine-tune CHOIS on the TRUMANS dataset.
@@
Quantitative evaluations cover three axes: object locomotion, human-scene interaction, hand-object interaction. The user study employs 30 sets of comparison videos assessing generation performance. Detailed metric explanation, visual comparisons, and more implementation details are provided in the supplementary material.
Ablation Study We validate our two key designs through controlled experiments:
  • Scene-Aware: Removes local scene information $B^{s}$ and relative scene distance loss function $\mathcal{L}_{sd}$.
  • Scene-Guided: Disables classifier-free guidance (Eq. 8), relying solely on joint-based optimization.
These experiments confirm the necessity of both local scene constraints and training-free scene guidance mechanisms. As for the efficiency of our navigation algorithm, it is demonstrated in the evaluation of object locomotion in Table 1.
@@
Figure 5. User Study. Users are asked to rate generated motions based on four different indicators, with the best being 3 points, the middle being 2 points, and the worst being 1 point. The scores are then tallied and the proportion of each method is calculated. The best result that could be obtained is 50%. The closer it is to 50%, the better the performance.
4.2. Results
4.2.1. Comparison
Quantitative Experiments Our method demonstrates three key advantages through comprehensive benchmarking. As shown in Table 1, HOSIG achieves a 93.3% success rate in object locomotion tasks, outperforming LINGO (23.3%) and CHOIS (3.3%) by factors of 4.0$\times$ and 28.3$\times$ respectively. This quantifies our method’s enhanced controllability in full-body motion generation, particularly in complex navigation-objective coordination scenarios. The hierarchical scene perception mechanism yields a penetration rate of 42% - 55% lower than baselines. While LINGO shows lower mean penetration volume (0.37 vs 0.66), our method prevents catastrophic failures as evidenced by maximum penetration values: 201.8 vs CHOIS’ 4,271.8. This confirms our layered constraint system effectively balances micro/macro scene interactions. As comparable contact rates (98% vs 24-28%), qualitative analysis reveals our full-body generation enables functional hand-object interactions. Although LINGO and CHOIS have low metrics on the penetration volume, this is mainly due to their extremely low hand-object contact rate. Overall, ours achieves better full-body HOSI generation.
The results validate our core hypotheses: 1) Hierarchical scene constraints (Pene. Rate improvement) enable collision-aware motion synthesis, 2) Full-body kinematics modeling (Contact. Rate stability) supports precise object manipulation, and 3) The dual guidance mechanism achieves optimal task-scene-object trade-offs (bold metrics).
@@ 4.2.1. Comparison
User Study The user study involved participants evaluating motion generation outcomes through four quality metrics: Motion Quality, Trajectory Quality, Manipulation Quality, and Overall Quality. Performance assessment is conducted by measuring the relative contribution of each method to the aggregated score, with metric definitions provided in Figure 7 and the Appendix. As depicted in Figure 5, our HOSIG method outperforms both baselines across all metrics, achieving values approaching 50% – the theoretical maximum for each indicator. This demonstrates that our method is consistently preferred over alternatives in pairwise comparisons. The subjective evaluations align with quantitative experimental results, closely matching the observed performance hierarchy HOSIG ¿ LINGO ¿ CHOIS. Complementary qualitative comparisons in Figure 6 reveal HOSIG’s consistent superiority across various scenarios. These findings collectively demonstrate that hierarchical scene perception enables HOSIG to achieve robust human-object-scene interaction synthesis.
4.2.2. Ablation Study
@@
Scene-Aware The Scene-Aware variant (removing local point cloud constraints) reveals critical insights into our spatial reasoning mechanism. As shown in Table 2, the full SGAP model improves scene interaction metrics by 16.3% (penetration rate), 69.9% (mean penetration volume), and 59.4% (max penetration volume) compared to the ablated version. Crucially, object interaction precision remained stable with ¡2% variation in grasp success rate, confirming that scene awareness enhances environment adaptation without compromising manipulation capabilities. Further analysis demonstrated that models trained without the scene distance loss (Eq. 2) suffer from higher penetration in scene avoidance, as unstructured scene features introduce noise in the latent space. In conclusion, SGAP not only needs additional scene information $B^{s}$ to perceive the scene, but also needs corresponding loss $\mathcal{L}_{sd}$ to learn how to use the scene information. Both scene information and scene distance loss are indispensable.
Table 2. Ablation studies of SGAP. The w/o sd means training SGAP without scene distance loss. The w/o sa means completely removing scene perception, including scene distance loss constraints and scene information $B^{s}$ input to the model.
@@
Table 3. Ablation studies of SCoMoGen. The w/o sg means inferencing without scene guidance $\mathcal{L_{s}}$.
@@
Scene-Guided The Scene-Guided configuration (disabling Eq. 8 in classifier-free guidance) exhibits significant performance degradation across all scene interaction metrics. As shown in Table 3, quantitative results showe higher penetration volumes and lower object locomotion stability compared to our full model. Notably, the guidance mechanism improves hand-object alignment precision. Experiments show that scenes are worth using as an additional modality to help generate motions. This is not only to avoid the penetration of human and scenes, but also to improve the quality of humans’ activities in the scene, such as manipulating objects.
4.3. Limitation and Future Work
While our framework demonstrates robust navigation and interaction capabilities, two limitations warrant discussion. First, the decoupled architecture introduces computational overhead that prevents real-time deployment. Second, the current implementation focuses on grasping, lacking fine-grained in-hand manipulation capabilities. Future work will pursue an end-to-end trainable system, which can autonomously find paths and generate motions in real-time. Additionally, integrating physics-based hand controllers with our scene-aware guidance could bridge the gap between global motion planning and dexterous object interaction.
5. Conclusions
@@
We present HOSIG, a hierarchical framework for synthesizing high-fidelity full-body human-object-scene interactions in complex 3D environments. By decoupling the task into scene-aware grasp pose generation, heuristic navigation planning, and scene-guided controllable motion synthesis, our method addresses critical limitations in existing HOI and HSI approaches, which often suffer from scene penetrations or fail to coordinate fine-grained manipulations with global navigation. The proposed hierarchical scene perception mechanism combines local geometry constraints, compressed 2D obstacle-aware maps, and dual-space diffusion guidance. It ensures collision-free interactions while maintaining precise hand-object contact and natural locomotion. Experiments demonstrate performance, with a 93.3% success rate in object locomotion tasks and significant reductions in scene penetration rates (42%–55%). Notably, our framework achieves unlimited motion length through autoregressive generation and requires minimal manual intervention, making it practical for applications in VR, robotics, and animation.
Appendix A Evaluation Metrics
A.1. Object Locomotion Assessment
The spatial-temporal performance of object manipulation is quantified through three core metrics.
Dist. The Distance metric measures the Euclidean distance between the object’s final position $p_{\text{final}}$ and the target location $p_{\text{target}}$, expressed as $\|p_{\text{final}}-p_{\text{target}}\|_{2}$ in meters. This metric directly reflects the system’s terminal positioning accuracy and spatial control capability, where smaller values indicate better alignment with the target configuration.
Time The Time metric captures the temporal efficiency of task execution, calculated as $t_{\text{end}}-t_{\text{start}}$ in seconds. It evaluates the overall planning optimality and motion coordination effectiveness, with lower values representing faster task completion without unnecessary detours or oscillations.
Rate The Success Rate provides a threshold-based binary evaluation, defined as the proportion of trials where the final distance falls within 0.05 meters: $\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}(\|p_{\text{final}}^{(t)}-p_{\text{target}}\|_{2}\leq 0.05)$. This metric serves as a comprehensive indicator of system reliability, combining spatial precision and temporal consistency requirements.
A.2. Human Interaction Analysis
To assess the physical plausibility of human-scene and human-object interactions, we introduce penetration metrics based on signed distance fields (SDF).
Pene. / Contact Rate The Penetration Rate quantifies the frequency of invalid collisions, computed as the temporal average of penetration occurrence:
(A1) $\mathcal{R}=\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}(\exists\mathbf{v}\in\mathcal{V}_{h}^{t}:\phi(\mathbf{v},M)<0)$
where $\mathcal{V}_{h}$ is the human vertices, $M$ is the scene/object mesh and $\phi(\cdot)$ is SDF function, $T$ is the total number of human poses. The Penetration Rate measure reveals the system’s ability to maintain basic collision avoidance constraints during dynamic interactions. As for Contact Rate, it evaluates the quality of hand-object interaction. With a higher contact rate, the object is less likely to be suspended in the air, which means that the motion of manipulating the object is more physically possible.
Pene. Mean The Mean Penetration Volume evaluates the spatial extent of penetration artifacts through depth-weighted integration:
(A2) $\mathcal{V}_{mean}=\frac{1}{T}\sum_{t=1}^{T}\sum_{\mathbf{v}\in\mathcal{V}_{h}^{t}}|\phi(\mathbf{v},M)|\cdot\mathbb{I}(\phi(\mathbf{v},O)<0)$
This metric characterizes the average penetration severity, with lower values indicating better contact modeling fidelity.
Pene. Max The Maximum Penetration Depth identifies worst-case penetration scenarios:
(A3) $\mathcal{V}_{max}=\max_{t\in[1,T]}\left[\sum_{\mathbf{v}\in\mathcal{V}_{h}^{t}}|\phi(\mathbf{v},M)|\cdot\mathbb{I}(\phi(\mathbf{v},O)<0)\right]$
This conservative measure highlights local failures in collision handling that may lead to visually implausible interactions, even when average penetration volumes appear acceptable.
A.3. User Study
Four dimensions are established for perceptual assessment:
Motion Quality evaluates the naturalness of human body movements in isolation. This metric focuses on biomechanical plausibility of limb motions and posture transitions during object manipulation, specifically assessing: (1) natural walking gaits, (2) physically valid grasping/releasing gestures, and (3) absence of unnatural body contortions. Scene context and object dynamics are visually masked during evaluation.
Trajectory Quality measures the spatial rationality of human navigation paths. Evaluators analyze 2D trajectory plots to assess: (1) path optimality (minimal detours), (2) motion smoothness (absence of abrupt turns), and (3) obstacle avoidance capability. Temporal aspects are excluded by using static trajectory visualizations.
Manipulation Quality focuses on human-object physical interaction validity. This criterion examines: (1) hand-object contact consistency during grasping, (2) object stability during transportation (no penetration or unrealistic oscillations), and (3) smooth object motion transitions. Finger-level details are ignored to ensure fair comparison, since two baseline methods can not generate hand poses.
Overall Quality provides integrated assessment of the complete motion synthesis system. Evaluators consider the synergistic effects of: (1) coordination between body motion and navigation trajectory, (2) physical plausibility of combined human-scene-object interactions, and (3) holistic visual realism of the generated motions.
Appendix B Implementation Details
This section delineates the critical implementation specifics that bridge our methodological framework with practical execution. The architecture exposition progresses through three coherent dimensions: (1) Systemic orchestration of the core computational pipeline, highlighting the synergistic interaction between principal components and stability-enhancing design choices; (2) Detailed exposition of training protocols encompassing parameter initialization strategies, optimization dynamics, and convergence monitoring mechanisms; (3) Comprehensive description of experimental configurations including dataset curation procedures, and benchmarking methodologies. Collectively, these implementation facets establish the technical foundation for reproducible empirical validation of our proposed framework.
B.1. Pipeline
B.1.1. Scene-Aware Grasp Pose Generation
The grasp pose generation module produces 10 candidate poses per object-scene configuration. To address implausible grasps caused by interpenetration between the character and the scene, we implement a collision-aware filtering mechanism. Each candidate pose undergoes collision detection against all scene elements. The pose with minimal aggregate collision volume is selected for subsequent processing. This mechanism effectively eliminates physically inconsistent grasps while maintaining sufficient pose diversity.
B.1.2. Heuristic Navigation Adjustment
For 2D obstacle-aware navigation, we enforce strict node validity checks during path planning. Any target node falling inside obstacle regions or causing pathfinding failures triggers an interactive adjustment protocol. The system visually overlays obstacle heatmaps on the navigation interface and allows human operators to reposition invalid nodes. This human-in-the-loop correction ensures all navigation targets reside in collision-free regions before executing path planning algorithms.
B.1.3. Motion Generation Constraints
The motion synthesis module employs three key constraints for temporal coherence:
1) Finger articulation constraints: Maintain fixed finger joint angles within 10 frames (5 frames before/after grasp events) to ensure stable object manipulation.
2) Motion continuity: Enforce 10-frame overlap between consecutive motion segments, using the final pose of the preceding motion as the initial condition for subsequent motion generation.
3) Object trajectory synchronization: Gradually transition object poses relative to the character hand during transportation phases. This smooth interpolation bridges the discrete “pick up” and “put down” configurations through spherical linear interpolation, preventing abrupt object position jumps.
B.2. Training Details
Our three-stage training pipeline utilizes the Trumans (Jiang et al., 2024b) dataset on a single NVIDIA RTX 3090/4090 GPU. Phase 1 trains the SGAP for 8 hours using the standard $\mathcal{L}_{VAE}$ loss and collision penalty terms. Following MDM, as a diffusion model, phase 2 trains the SCoMoGen’s main branch over 48 hours.
Phase 3 strictly follows ControlNet’s paradigm: 1) Duplicate main branch weights to initialize the control branch. 2) Zero-initialize trainable control layers. 3) Freeze main network while training only control-specific parameters through 24-hour fine-tuning. This strategy injects spatial control signals via residual connections, optimized with $\mathcal{L}_{diff}$ plus $\mathcal{L}_{ctrl}=\|c_{target}-c_{pred}\|_{2}$ where $c$ denotes control embeddings. Our implementation builds on these open-source works (Taheri et al., 2022; Li et al., 2024a; Xie et al., 2023). And we will also make our code public as soon as possible after publication.
B.3. Experiment Setup
Evaluation of HOSI Generation The evaluation protocol assesses two critical capabilities: object-agnostic grasping proficiency and scene generalization. We select three daily objects (bottle, cup, mouse) and 10 scenes excluded from the training set. Each scene-object pair contains randomized initial and target SE(3) poses, creating 30 out-of-distribution test cases. These configurations strictly avoid spatial overlaps with training data to validate cross-scene robustness.
Evaluation of Grasp Pose Generation As shown in Table 2 in the paper, for the ablation study on grasp generation, SGAP generates 10 grasp candidates per object pose using identical test conditions, resulting in 600 evaluation trials (30 cases × 10 attempts).

### v2

@@ Introduction
Ours 0.0270 13.0367 0.9333 0.1851 0.6562 201.8264 0.9800 0.0007 0.0113
Table 1: Quantitative results of human-object-scene interaction generation. This mainly involves the interaction between human and scenes, human and objects when characters operate objects in the scene. At the same time, this also evaluates the efficiency and accuracy of the character in carrying objects.
@@
$\mu_{t}=\mu_{t}-\tau\nabla_{\mu_{t}}\left(\lambda_{1}\mathcal{L}_{j}^{root}+\lambda_{2}\mathcal{L}_{j}^{hand}+\lambda_{3}\mathcal{L}_{s}\right)$ (6)
@@
Experiments
Implementation Details
@@
Our evaluation protocol comprehensively assesses both absolute performance and component effectiveness. All experiments are conducted on the TRUMANS dataset (Jiang et al. 2024b), containing 100 indoor scenes with annotated human-object interactions. We establish two evaluation axes: 1) comparison against SOTA methods in HOI and HSI generation, and 2) ablation studies isolating our core technical innovations.
Metrics We employ two complementary metric categories: Object Locomotion Assessment evaluates spatial-temporal performance through: (1) terminal positioning accuracy (Dist: Euclidean distance to target), (2) temporal efficiency (Time: task duration), and (3) reliability (Rate: success proportion within 0.05m threshold). Human Interaction Analysis quantifies physical plausibility via SDF-based penetration metrics: collision frequency (Penetration Rate), hand-object contact quality (Contact Rate), average severity (Mean Penetration Volume), and worst-case failures (Max Penetration Depth).
@@
SOTA Comparisons While no existing method achieves identical functionality, we select two representative baselines LINGO (Jiang et al. 2024a) and CHOIS (Shi et al. 2023) for fair comparison. Quantitative evaluations cover three axes: object locomotion, human-scene interaction, hand-object interaction. The user study employs 30 sets of comparison videos assessing generation performance. Detailed metric explanation, visual comparisons, and more implementation details like baseline reproduction and training details, are provided in the supplementary material.
Ablation Study We validate our three key designs through controlled experiments: (1) Scene-Aware: Removes local scene information $B^{s}$ and relative scene distance loss function $\mathcal{L}_{sd}$. (2) Heuristic Navigation: No additional experiments were set up, and its efficiency can be reflected in Table 1. (3) Scene-Guided: Disables gradient-based guidance (Eq. 5), relying solely on joint-based optimization.
Results
Comparison
[…]
@@ Results
Figure 6: User Study. Users are asked to rate generated motions based on four different indicators, with the best being 3 points, the middle being 2 points, and the worst being 1 point. The scores are then tallied and the proportion of each method is calculated. The best result that could be obtained is 0.50. The closer it is to 0.50, the better the performance.
@@
User Study The user study involved participants evaluating motion generation outcomes through four quality metrics: Motion Quality, Trajectory Quality, Manipulation Quality, and Overall Quality. Performance assessment is conducted by measuring the relative contribution of each method to the aggregated score, with metric definitions provided in supplementary materials. As depicted in Figure 6, our HOSIG method outperforms both baselines across all metrics, achieving values approaching 50% – the theoretical maximum for each indicator. This demonstrates that our method is consistently preferred over alternatives in pairwise comparisons. The subjective evaluations align with quantitative experimental results, closely matching the observed performance hierarchy HOSIG >> LINGO >> CHOIS. These findings collectively demonstrate that hierarchical scene perception enables HOSIG to achieve robust human-object-scene interaction synthesis.
Ablation Study
@@
Scene-Aware The Scene-Aware variant (removing local point cloud constraints) reveals critical insights into our spatial reasoning mechanism. As shown in Table 2, the full SGAP model improves scene interaction metrics by 16.3% (penetration rate), 69.9% (mean penetration volume), and 59.4% (max penetration volume) compared to the ablated version. Crucially, object interaction precision remained stable with $2\%$ variation in grasp success rate, confirming that scene awareness enhances environment adaptation without compromising manipulation capabilities. Further analysis demonstrated that models trained without the scene distance loss (Eq. 2) suffer from higher penetration in scene avoidance, as unstructured scene features introduce noise in the latent space. In conclusion, SGAP not only needs additional scene information $B^{s}$ to perceive the scene, but also needs corresponding loss $\mathcal{L}_{sd}$ to learn how to use the scene information. Both scene information and scene distance loss are indispensable.
Heuristic Navigation The efficacy of our Heuristic Navigation framework manifests most prominently in two critical metrics: temporal efficiency (Time in Object Locomotion) and collision integrity (Penetration Rate in Scene Interaction). As quantified in Table 1, our approach achieves a $3.6$ second average improvement in task completion time over the LINGO baseline. This significant acceleration stems directly from global scene comprehension, which enables anticipatory obstacle avoidance and optimal route planning. Conversely, local perception methods frequently encounter pathological scenarios, particularly in geometrically complex environments. Myopic decision making leads to navigation dead-ends and recovery behaviors, as qualitatively demonstrated in our supplementary visualizations. Furthermore, our method’s superior Penetration Rate ($\downarrow 23\%$ versus LINGO) also confirms enhanced path precision.
@@
Table 2: Ablation studies of SGAP. The w/o sd means training SGAP without scene distance loss. The w/o sa means completely removing scene perception, including scene distance loss constraints and scene information $B^{s}$ input to the model.
@@
Table 3: Ablation studies of SCoMoGen. The w/o sg means inferencing without scene guidance $\mathcal{L_{s}}$.
@@
Scene-Guided The Scene-Guided configuration (disabling Eq. 5 in gradient-based guidance) exhibits significant performance degradation across all scene interaction metrics. As shown in Table 3, quantitative results show higher penetration volumes and lower object locomotion stability compared to our full model. Notably, the guidance mechanism improves hand-object alignment precision. Experiments show that scenes are worth using as an additional modality to help generate motions. This is not only to avoid the penetration of human and scenes, but also to improve the quality of humans’ activities in the scene, such as manipulating objects.
Conclusions
@@
We present HOSIG, a hierarchical framework for synthesizing high-fidelity full-body human-object-scene interactions in complex 3D environments. By decoupling the task into scene-aware grasp pose generation, heuristic navigation planning, and scene-guided controllable motion synthesis, our method addresses critical limitations in existing HOI and HSI approaches. The proposed hierarchical scene perception mechanism combines local geometry constraints, compressed 2D obstacle-aware maps, and dual-space diffusion guidance. It ensures collision-free interactions while maintaining precise hand-object contact and natural locomotion. Notably, our framework achieves unlimited motion length through autoregressive generation and requires minimal manual intervention, making it practical for applications in VR, robotics, and animation.